OpenStudy (anonymous):

50% less than greater than or equal to 2/7

OpenStudy (anonymous):

50% = 50/100 = 1/2 1/2 is greater than 2/7. Why? because half of 7 is 3.5 so it needs to be 4/7 or 5/7 or 6/7 to be greater than 1/2